Find the critical value z Subscript alpha divided by 2 that corresponds to alpha = 0.19.
Answer #1

Solution:

Given that,

= 0.19

/ 2 = 0.19 / 2 = 0.095

Z/2 = Z0.095 = 1.311

critical value = 1.311